the current average of monthly salary of the new school public teacher is 23,250. The net monthly salary is nominal distributed with a standard deviation of 5,725


Find the probability that a teacher receives a net monthly income of 25,000


1
Expert's answer
2021-11-16T06:05:31-0500

"P(x=25000)=P(24995<x<25005)=P(x<25005)-P(x<24995)"


"z_2=\frac{25005-23250}{5725}=0.31"


"z_1=\frac{24995-23250}{5725}=0.30"


"P(24995<x<25005)=P(z<0.31)-P(z<0.30)=0.6217-0.6179=0.038"